OpenStudy (anonymous):

The equation sec^(2) x -1 = tan^2 x is an idenity. True or false.

OpenStudy (anonymous):

sec^2(x) = 1/cos^2(x) so your equation is : [1/cos^2(x)] - 1 = tan^2(x) multiply both sides by co^2(x) we get 1 - cos^2(x) = sin^2(x) 1 = cos^2(x) + sin^2(x) which is a known identity

OpenStudy (anonymous):

in fact sec^(2) x -1 = tan^2 x or more commonly tan^2(x) + 1= sec^2(x) is also a known identity :)

OpenStudy (anonymous):

\[x ^{2}+y ^{2}=r ^{2}\]If you have the pythagorean theorem: Then try to get tangent and see what it equals. Because tangent is y/x i'll try to put it in this form... So you can see the identity for tangent.